# A Recursive Theory of Variational State Estimation: The Dynamic Programming Approach

### v1

@@ Abstract
In this article, variational state estimation is examined from the dynamic programming perspective. This leads to two different value functional recursions depending on whether backward or forward dynamic programming is employed. The result is a theory of variational state estimation that corresponds to the classical theory of Bayesian state estimation. More specifically, in the backward method, the value functional corresponds to a likelihood that is upper bounded by the state likelihood from the Bayesian backward recursion. In the forward method, the value functional corresponds to an unnormalized density that is upper bounded by the unnormalized filtering density. Both methods can be combined to arrive at a variational two-filter formula. Additionally, it is noted that optimal variational filtering is generally of quadratic time-complexity in the sequence length. This motivates the notion of sub-optimal variational filtering, which also lower bounds the evidence but is of linear time-complexity. Another problem is the fact that the value functional recursions are generally intractable. This is briefly discussed and a simple approximation is suggested that retrieves the filter proposed by Courts2021. The methodology is examined in a jump Gauss–Markov system, where it is observed that the value functional recursions are tractable under a certain factored Markov process approximation. A simulation study demonstrates that the posterior approximation is of adequate quality.
@@ 1.1 Bayesian state estimation theory
There are essentially two ways to represent the a posteriori distribution over state paths, the forward recursion and the backward recursion (Cappe2005), see also Ho1964, Askar1981, Kitagawa1987. The result in a forward and reverse-time Markov factorization of the posterior, respectively. In order to state the result, introduce the following notation. The conditional distribution of $x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}$ given $x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}$ and $y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}$ be denoted by
@@
$\bar{\pi}_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}\mid t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}^{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}(x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}\mid x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}},y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}})=h_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}\mid t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}(y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}\mid x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}})\pi_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}\mid t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}^{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}(x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}\mid x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}),$ (4)
where the ranges $t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}$ and $t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}$ are non-overlapping. For example, $\bar{\pi}_{t\mid t-1}^{t}=h_{t\mid t}\pi_{t\mid t-1}$, the unnormalized filtering density satisfy $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=h_{0\mathrel{\mathop{\ordinarycolon}}t}\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$, and the unnormalized smoothing density satisfy $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}=h_{0\mathrel{\mathop{\ordinarycolon}}T}\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$. The backward method establishes a recursion over $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ from which a forward Markov representation of $\pi_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$ is obtained, this is theorem˜1.
@@ 1.2 Contribution and outline
Recursive approaches have previously been formulated by Weber2015 and Campbell2021 from the reinforcement learning perspective. Here an optimal control approach is taken instead where the time-marginals of $q_{0\mathrel{\mathop{\ordinarycolon}}T}$ are states and the associated transition kernels are the controls. The problem can then formally be solved by both backward and forward dynamic programming (Bellman1966, Larson1966), which involve recursions over associated value functionals.  [^1] It ought to be noted that the present notion of value functionals differs from the notion of value function used by Weber2015, Campbell2021. These methods shall be referred to as the variational backward and forward methods, respectively. Based on this a theory of recursive variational state estimation is developed in section˜2, which goes beyond the contributions of Weber2015, Campbell2021.  [^2] The relationship between the present contribution and that of Weber2015, Campbell2021 is further discussed in section 5. More specifically, a variational analogue of the Bayesian state estimation is obtained by establishing:
@@
The recursions of theorem˜3 and theorem˜2 are parametrized by the time-marginals of the variational posterior. This problem is circumvented by the fixed-point equations of propositions˜1 and 2. Furthermore, the notion of variational filtering is discussed in section˜3. In particular, the preceeding analysis suggests that the optimal variational filter is in general of time-complexity $O(T^{2})$. This leads to the notion of sub-optimal variational filtering, which is of time-compelxity $O(T)$. Another problem is that the value functional recursions are intractable in general. This is briefly discussed in section˜4, where the fact that $\alpha_{t}$ is an unnormalized density is used together with variational inference to obtain an approximation. This procedure is shown to recover the filter proposed by Courts2021. Additionally, in section˜6 the methodology is studied in a jump-Gauss Markov systems when the a posterior distribution of the state and jump processes is approximated as independendent Markov processes. In this case the value functional recursions are tractable. The sub-optimal filter and the fixed-point iterations are studied in simulations and are found to give reasonable results.
@@ 2.1 Markov Factorization of the variational family
$q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})=q_{T}(x_{T})\prod_{t=1}^{T}q_{t\mid t+1}(x_{t}\mid x_{t+1}).$ (18)
@@
In section˜2.2, the forward Markov factorization (16) is employed together with backward dynamic programming, giving a variational anlogue of theorem˜1. In section˜2.3 the reverse-time Markov factorization (18) is employed together with forward dynamic programming, which gives a variational analogue of theorem˜2. Lastly, in section˜2.4 the mixed forward/reverse-time factorization (19) is employed together with both forward and backward dynamic programming to obtain a variational analogue of corollary˜1.
@@ 2.2 Forward Markov factorization of the variational family
The boundary constraint $V^{\mathsf{b}}_{T}(q_{T})=0$ of definition˜3 ensure that this recursion is valid for $0<t\leq T$. It is important to note that $q_{t}$ was obtained by marginalizing $q_{t\mid t-1}$ with respect to $q_{t-1}$. Respecting this constraint gives the following result.
@@ 3 Optimal and sub-optimal variational filtering
In many applications it is interesting to compute an approximation of the filtering density, $\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$, on-line. Now in view of theorem˜4 and corollary˜3, the variational filtering densities may be defined as follows.
@@
$q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=\underset{}{\arg\,\max}\,\int\log\Big\{\frac{\alpha_{t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}(x_{t})\operatorname{d\!}x_{t}.$ (46)
The variational predictive densities, $q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t-1}$, are defined similarly by replacing $\alpha_{t}$ with $\alpha_{t}^{\mathsf{p}}$.
A problem of the optimal filter given by definition˜5 is that it will in general be of time-complexity $O(T^{2})$ (c.f proposition˜2). More specifically, in general the sequence $\alpha_{1},\ldots,\alpha_{t}$ depend on the variational filtering density at time $t$, via the past time-marginals (c.f. section˜2.3).  [^3] In principle this would be clearer with the notation $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}s}$ for forward value functionals based on observations $y_{0\mathrel{\mathop{\ordinarycolon}}s}$. Consequently, the optimal variational filter can not be found without back-tracking, which entail a time-complexity $O(T^{2})$.
However, there is a special case when the optimal variational filter is of time-complexity $O(T)$. Define the universally optimal predictive value functional, $\alpha_{t}^{\dagger}$, and the universally optimal reverse-time transition kernel, $q_{t-1\mid t}^{\dagger}$ with respect to $\alpha_{t-1}$, by Bayes’ rule
@@
$\begin{split}&\int\log\{\alpha_{t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}=\int\log\{h_{t\mid t}(x_{t})\alpha_{t}^{\dagger}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}\\ &+\underset{q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\max}\,\int\Big(\int\log\Big\{\frac{q_{t-1\mid t}^{\dagger}(x_{t-1}\mid x_{t})}{q_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}q_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\Big)q_{t}(x_{t})\operatorname{d\!}x_{t}.\end{split}$
@@
Assume that $q_{t-1\mid t}^{\dagger}\in\mathcal{Q}_{t-1\mid t}$, then the forward value functional satisfies the following recursion
$\log\alpha_{t}(x_{t})=\log\{h_{t\mid t}(x_{t})\alpha_{t}^{\dagger}(x_{t})\}.$
@@
$\alpha_{0}$ $=\bar{\pi}_{0}^{0}$ (48a)
$q_{0}^{0}$ $=\underset{q_{0}\in\mathcal{Q}_{0}}{\arg\,\max}\,\int\log\Big\{\frac{\alpha_{0}}{q_{0}}\Big\}q_{0}\operatorname{d\!}x_{0}$ (48b)
$q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t},\,q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}$ $=\underset{q_{t}\in\mathcal{Q}_{t},\,q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\arg\,\max}\,\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}}{q_{t-1\mid t}q_{t}}\Big\}q_{t-1\mid t}q_{t}\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t}$ (48c)
$\log\alpha_{t}$ $=\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}}{q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}}\Big\}q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}\operatorname{d\!}x_{t-1}.$ (48d)
While the procedure of definition˜6 is indeed sub-optimal, it can be seen that $\hat{\alpha}_{t}$ in fact also provides a lower bound on the unnormalized filtering density $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$, just like $\alpha_{t}$. It is just not the optimal bound obtainable with respect to the chosen variational family. Furthermore, the sub-optimal variational filter serves as a reasonable method of initializing the fixed-point iteration of proposition˜2.

### v2

@@ Abstract
In this article, variational state estimation is examined from the dynamic programming perspective. This leads to two different value functional recursions depending on whether backward or forward dynamic programming is employed. The result is a theory of variational state estimation that corresponds to the classical theory of Bayesian state estimation. More specifically, in the backward method, the value functional corresponds to a likelihood that is upper bounded by the state likelihood  [^1] This is often referred to as the “backward filter”. from the Bayesian backward recursion. In the forward method, the value functional corresponds to an unnormalized density that is upper bounded by the unnormalized filtering density. Both methods can be combined to arrive at a variational two-filter formula. Additionally, it is noted that optimal variational filtering is generally of quadratic time-complexity in the sequence length. This motivates the notion of sub-optimal variational filtering, which also lower bounds the evidence but is of linear time-complexity. Another problem is the fact that the value functional recursions are generally intractable. This is briefly discussed and a simple approximation is suggested that retrieves the filter proposed by Courts2021. The methodology is examined in (i) a jump Gauss–Markov system under a certain factored Markov process approximation, and (ii) in a Gauss–Markov model with log-polynomial likelihoods under a Gauss–Markov constraint on the variational approximation. It is demonstrated that the value functional recursions are tractable in both cases. The resulting estimators are examined in simulation studies and are found to be of adequate quality in comparison to sensible baselines.
@@ 1.1 Bayesian state estimation theory
There are essentially two ways to construct the a posteriori distribution over state paths, the forward recursion and the backward recursion (Cappe2005), see also Ho1964, Askar1981, Kitagawa1987. The result in a forward and reverse-time Markov factorization of the posterior, respectively. In order to state the result, introduce the following notation. The conditional distribution of $x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}$ given $x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}$ and $y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}$ be denoted by
@@
$\bar{\pi}_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}\mid t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}^{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}(x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}\mid x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}},y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}})=h_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}\mid t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}(y_{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}\mid x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}})\pi_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}\mid t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}^{t_{5}\mathrel{\mathop{\ordinarycolon}}t_{6}}(x_{t_{1}\mathrel{\mathop{\ordinarycolon}}t_{2}}\mid x_{t_{3}\mathrel{\mathop{\ordinarycolon}}t_{4}}).$ (4)
For example, $\bar{\pi}_{t\mid t-1}^{t}=h_{t\mid t}\pi_{t\mid t-1}$, the unnormalized filtering density satisfy $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=h_{0\mathrel{\mathop{\ordinarycolon}}t}\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$, and the unnormalized smoothing density satisfy $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}=h_{0\mathrel{\mathop{\ordinarycolon}}T}\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$. The backward method establishes a recursion over $h_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}$ from which a forward Markov representation of $\pi_{0\mathrel{\mathop{\ordinarycolon}}T}^{0\mathrel{\mathop{\ordinarycolon}}T}$ is obtained, this is theorem˜1.
@@ 1.2 Contribution and outline
Recursive approaches have previously been formulated by Weber2015 and Campbell2021 from the reinforcement learning perspective. Here an optimal control approach is taken instead where the time-marginals of $q_{0\mathrel{\mathop{\ordinarycolon}}T}$ are states and the associated transition kernels are the controls. The problem can then formally be solved by both backward and forward dynamic programming (Bellman1966, Larson1966), which involve recursions over associated value functionals.  [^2] It ought to be noted that the present notion of value functionals differs from the notion of value function used by Weber2015, Campbell2021. The relationship between the present contribution and that of Weber2015, Campbell2021 is further discussed in section 5. These methods shall be referred to as the variational backward and forward methods, respectively. Based on this a theory of recursive variational state estimation is developed in section˜2, which goes beyond the contributions of Weber2015, Campbell2021.
More specifically, a variational analogue of the Bayesian state estimation is obtained by establishing:
@@
The recursions of theorem˜3 and theorem˜2 are parametrized by the time-marginals of the variational posterior, which are of course unavailable. This problem is circumvented by the fixed-point equations of propositions˜1 and 2. Furthermore, the notion of variational filtering is discussed in section˜3. In particular, the preceeding analysis suggests that the optimal variational filter is in general of time-complexity $O(T^{2})$. This leads to the notion of sub-optimal variational filtering, which is of time-compelxity $O(T)$. Another problem is that the value functional recursions are intractable in general. This is briefly discussed in section˜4, where the fact that $\alpha_{t}$ is an unnormalized density is used together with variational inference to obtain an approximation. This procedure is shown to recover the filter proposed by Courts2021.
Additionally, two case studies are given where the value functional recursions are indeed tractable. In section˜6 jump-Gauss Markov systems are studied when the a posterior distribution of the state and jump processes is approximated as independendent Markov processes and in section˜7 a Gauss–Markov model with log-polynomial likelihoods is examined under a Gauss–Markov constraint.
@@ 2.1 Markov Factorization of the variational family
$q_{0\mathrel{\mathop{\ordinarycolon}}T}(x_{0\mathrel{\mathop{\ordinarycolon}}T})=q_{T}(x_{T})\prod_{t=0}^{T-1}q_{t\mid t+1}(x_{t}\mid x_{t+1}).$ (18)
@@
where the fact that $\mathcal{V}(\bar{\pi}_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t}^{t+1\mathrel{\mathop{\ordinarycolon}}T},q_{t+1\mathrel{\mathop{\ordinarycolon}}T\mid t})$ is a function in $x_{t}$ is omitted from the notation. In section˜2.2, the forward Markov factorization (16) is employed together with backward dynamic programming, giving a variational anlogue of theorem˜1. In section˜2.3 the reverse-time Markov factorization (18) is employed together with forward dynamic programming, which gives a variational analogue of theorem˜2. Lastly, in section˜2.4 the mixed forward/reverse-time factorization (19) is employed together with both forward and backward dynamic programming to obtain a variational analogue of corollary˜1.
@@ 2.2 Forward Markov factorization of the variational family
The boundary constraint $V^{\mathsf{b}}_{T}(q_{T})=0$ of definition˜3 ensure that this recursion is valid for $0<t\leq T$. It is important to note that $q_{t}$ is obtained by marginalizing $q_{t\mid t-1}$ with respect to $q_{t-1}$. Respecting this constraint gives the following result.
@@ 3 Optimal and sub-optimal variational filtering
In many applications it is interesting to compute an approximation of the filtering density, $\pi_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$, on-line. Note that the dependence of the value functionals with respect to the observation sequence has hitherto been supressed from the notation. However, in order to appropriately discuss the possibility of variational filtering it needs to be there. That is, let $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}s}$ denote the forward value functional with respect to the observations $y_{\tau},0\leq\tau\leq s$. Now in view of theorem˜4 and corollary˜3, the variational filtering densities may be defined as follows.
@@
$q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=\underset{}{\arg\,\max}\,\int\log\Big\{\frac{\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})}{q_{t}(x_{t})}\Big\}q_{t}(x_{t})\operatorname{d\!}x_{t}.$ (46)
The problem is now immediately clear, theorem˜4 gives a recursion in $t$ of $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$. However, in order to obtain a recursive variational filter what is needed is a recursion of $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ in $t$. The present analysis suggests that this is in general impossible without re-processing previous observations. This indicates that the optimal filter given by definition˜5 is of time-complexity $O(T^{2})$ (c.f proposition˜2).
However, there is a special case when the optimal variational filter is of time-complexity $O(T)$. Define the universally optimal predictive value functional, $\alpha_{t}^{\dagger}$, and the universally optimal reverse-time transition kernel, $q_{t-1\mid t}^{\dagger}$ with respect to $\alpha_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}$, by Bayes’ rule
@@
$\begin{split}&\int\log\{\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}=\int\log\{h_{t\mid t}(x_{t})\alpha_{t}^{\dagger}(x_{t})\}q_{t}(x_{t})\operatorname{d\!}x_{t}\\ &+\underset{q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\max}\,\int\Big(\int\log\Big\{\frac{q_{t-1\mid t}^{\dagger}(x_{t-1}\mid x_{t})}{q_{t-1\mid t}(x_{t-1}\mid x_{t})}\Big\}q_{t-1\mid t}(x_{t-1}\mid x_{t})\operatorname{d\!}x_{t-1}\Big)q_{t}(x_{t})\operatorname{d\!}x_{t}.\end{split}$
@@
Assume that $q_{t-1\mid t}^{\dagger}\in\mathcal{Q}_{t-1\mid t}$, then the forward value functional satisfies the following recursion for $1\leq t\leq T$
$\alpha_{t}^{\dagger}(x_{t})$ $=\int\pi_{t\mid t-1}(x_{t}\mid x_{t-1})\alpha_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}(x_{t-1})\operatorname{d\!}x_{t-1}$ (48a)
$\log\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}(x_{t})$ $=\log\{h_{t\mid t}(x_{t})\alpha_{t}^{\dagger}(x_{t})\}.$ (48b)
Additionally, $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}=\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}T}$.
@@
$\alpha_{0}^{0}$ $=\bar{\pi}_{0}^{0}$ (49a)
$q_{0}^{0}$ $=\underset{q_{0}\in\mathcal{Q}_{0}}{\arg\,\max}\,\int\log\Big\{\frac{\alpha_{0}^{0}}{q_{0}}\Big\}q_{0}\operatorname{d\!}x_{0}$ (49b)
$q_{t}^{0\mathrel{\mathop{\ordinarycolon}}t},\,q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}$ $=\underset{q_{t}\in\mathcal{Q}_{t},\,q_{t-1\mid t}\in\mathcal{Q}_{t-1\mid t}}{\arg\,\max}\,\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}}{q_{t-1\mid t}q_{t}}\Big\}q_{t-1\mid t}q_{t}\operatorname{d\!}x_{t-1\mathrel{\mathop{\ordinarycolon}}t}$ (49c)
$\log\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ $=\int\log\Big\{\frac{\bar{\pi}_{t\mid t-1}^{t}\alpha_{t-1}^{0\mathrel{\mathop{\ordinarycolon}}t-1}}{q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}}\Big\}q_{t-1\mid t}^{0\mathrel{\mathop{\ordinarycolon}}T}\operatorname{d\!}x_{t-1}.$ (49d)
The procedure of definition˜6 is indeed sub-optimal but new definition of $\alpha_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ in fact also provides a lower bound on the unnormalized filtering density $\bar{\pi}_{t}^{0\mathrel{\mathop{\ordinarycolon}}t}$ – the argument is exactly the same as for corollary˜3. It is just not necessarily the optimal bound obtainable with respect to the chosen variational family. Furthermore, the sub-optimal variational filter serves as a reasonable method of initializing the fixed-point iteration of proposition˜2.